OpenStudy (anonymous):

What are the discontinuities of the function f(x) =x^2-16/4x +24

OpenStudy (anonymous):

What are the point(s) where the denominator equals 0? In other words, what is the solution to 4x + 24 = 0?

OpenStudy (anonymous):

x= -24/4?

OpenStudy (anonymous):

Yes, or in other words, -6.

OpenStudy (anonymous):

is that the discontinuity?

OpenStudy (anonymous):

That is a discontinuity.

OpenStudy (anonymous):

And there's your answer.

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

\(\bf f(x)=\cfrac{x^2-16}{{\color{brown}{ 4x +24}}}\qquad {\color{brown}{ 4x +24}}=0\qquad \cfrac{x^2-16}{{\color{brown}{ 0}}}\to undefined\)